The area of the piece is 30 ft². The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth of a foot.
Answer by nerdybill(7384)

The area of the piece is 30 ft². The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth of a foot.
.
Let w = width
then
3w-1 = length
.
w(3w-1) = 30
3w^2-w = 30
3w^2-w-30 = 0
replace middle term:
3w^2 +9w-10w -30 = 0
group terms:
(3w^2+9w) - (10w+30) = 0
3w(w+3) - 10(w+3) = 0
(w+3)(3w-10) = 0
w = {-3, 10/3}
throw out the negative solution (extraneous) leaving:
w = 10/3
w = 3.33 feet